OpenStudy (anonymous):

A lawn can be mowed at an average rate of 44 square feet per minute. Which is closest to this rate is square yards per second? A 0.01 B 0.08 C 0.24 D 0.31 Help please????

OpenStudy (anonymous):

B

OpenStudy (anonymous):

1 minute = 60 seconds (44/9)/60 44/9 **1/60 =44/540 =0.08sq.yard/second

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np